### v2

# Routing Guidance for Emerging Transportation Systems with Improved Dynamic Trip Equity

Abstract
In this paper, we present a dynamic routing guidance system that optimizes route recommendations for individual vehicles within an emerging transportation system while enhancing travelers’ trip equity. We develop a framework to quantify trip quality and equity in a dynamic travel environment, providing new insights into how routing guidance influences equity in road transportation. Our approach enables real-time routing by incorporating both monitored and anticipated traffic congestion. We provide conditions that ensure achieving perfect trip equity for all travelers in a free-flow network. Finally, simulation studies on 1,000 vehicles traversing an urban road network in Boston demonstrate that our proposed method improves trip equity by approximately 11.4% compared to the shortest-route strategy. In addition, the results reveal that our approach redistributes travel costs across vehicle types through route optimization, contributing to a more equitable transportation system.
I Introduction
Transportation is a fundamental pillar of modern society, enabling access to essential daily services and activities, such as education, employment, healthcare, and social interactions. Recent advancements in transportation technologies, including shared mobility [1], electric vehicles [2, 3], and automated driving systems [4], are transforming travel patterns. Specifically, shared mobility reduces reliance on private vehicle ownership, electric vehicles lower travel costs while mitigating greenhouse gas emissions, and automated driving enhances road safety and traffic efficiency. These innovations hold great potential to shape a more efficient, cost-effective, and sustainable transportation system.
Equity has been an important principle in transportation to promote fairness and inclusivity [5]. It advocates for the just distribution of transportation resources, services, and infrastructure, ensuring that all individuals, regardless of their socioeconomic status, race, gender, or geographic location, have equal access to safe, affordable, and reliable mobility options. Transportation solutions lacking access equity can lead to unintended negative consequences [6]. As noted in [7], low-income populations, who often rely on public transit, face substantial barriers when affordable and reliable transportation options are limited. This inequity restricts their access to essential services and opportunities and reinforces cycles of poverty and social exclusion. In addition, disparities in transportation access can drive over-reliance on private vehicles, exacerbating traffic congestion, increasing travel costs, and contributing to environmental degradation.
Advancing equity in emerging transportation systems requires developing routing guidance systems that address the diverse needs of travelers, such as availability and affordability, while prioritizing equity. Different from conventional routing guidance services [8], which focus on optimizing traffic efficiency through metrics such as shortest distance, minimal time, or cost, equity-based routing guidance systems integrate fairness into existing frameworks, delivering socially optimal solutions for every traveler despite their disparities.
Existing routing guidance systems fall into two categories: static and dynamic systems [9]. Static routing guidance systems (S-RGSs) rely on macroscopic approaches to generate flow-level route recommendations, offering uniform guidance to all vehicles in the same flow. These systems are less accurate in computing optimal routes as they fail to incorporate real-time traffic information, such as congestion or incidents, for updating routes. Moreover, since all vehicles in the same flow are treated the same, static routing strategies may shift congestion from one point to another rather than effectively resolve it [10]. In contrast, dynamic routing guidance systems (D-RGSs) leverage real-time traffic data to continuously update route recommendations, resulting in more accurate and adaptive routes [11, 12]. While dynamic systems offer improved accuracy, studies on developing effective D-RGSs are still scarce, particularly in the context of enhancing mobility equity.
In this paper, we present a D-RGS designed to optimize the route recommendations for individual vehicles in an emerging transportation system with enhanced trip equity. Unlike existing D-RGSs, our approach integrates equity considerations into the decision-making process, ensuring that the benefits of transportation advancements are distributed fairly across diverse populations. By leveraging emerging technologies, including automated and ride-hailing vehicles, our system aims to bridge the gap between socioeconomic groups, reduce racial and ethnic disparities, and promote sustainable transportation alternatives. The contributions of this work are summarized as follows:
  • We present a novel framework to measure the quality and equity of trips in a dynamic emerging transportation system through the notions of the dynamic trip index and dynamic trip equity.
  • Based on the proposed metrics, we derive conditions for achieving perfect trip equity in free-flow networks, which provides the upper bound for trip equity attainable through route optimization.
  • We propose a D-RGS that incorporates both monitored and estimated traffic congestion in real-time route optimization and allows for trip equity improvement.
Finally, simulation studies on $1,000$ vehicles traversing an urban road network in Boston demonstrate the effectiveness of the proposed framework.
II Problem Formulation
In this section, we formulate the routing problem for individual vehicles. We start by introducing the road network, vehicle dynamics, and traffic flow. Subsequently, we formally present the routing problem.
II-A Road Network
We model the road network as a directed graph $\mathcal{G}(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes and $\mathcal{E}\!\subset\!{\mathcal{V}\!\times\!{\mathcal{V}}}$ is the set of edges. Each node $v\!\in\!\mathcal{V}$ is uniquely indexed, representing a key intersection or an origin/destination node in the network, while each edge $(v,v^{\prime})\!\in\!\mathcal{E}$ represents a road segment connecting a pair of nodes. The weight of each edge, denoted by $\tau(v,v^{\prime})\!\in\!\mathbb{R}_{+}$, denotes the travel time to traverse the edge, where $\mathbb{R}_{+}$ is the set of positive real numbers. Throughout this paper, we assume all vehicles travel with a constant free-flow speed. We denote the nominal (i.e., free-flow) travel time on an edge $(v,v^{\prime})$ by $\tau_{0}(v,v^{\prime})$.
Consider an emerging transportation system consisting of diverse transportation modes, including private vehicles, autonomous vehicles, and ride-hailing vehicles. At any time $t$, let $N(t)\!\in\!{\mathbb{N}}$ be the number of vehicles actively traveling in the network. By “actively,” we refer to the vehicle taking at least one traveler en route to a designated destination. For any vehicle $i\!\in\!\mathcal{N}(t):=\{1,2,\dots,N(t)\}$ traveling from a node $v_{1}$ to node $v_{n}$, a feasible route between the two nodes is denoted as
$r_{1,n}\!:=\!\big{\{}(v_{1},v_{2}),(v_{2},v_{3}),\dots,(v_{n-1},v_{n})\big{\}},$ (1)
where $(v_{k},v_{k+1})\in\mathcal{E}$, $k\!=\!1,\dots,n\!-\!1$. The set of all feasible routes from $v_{1}$ to $v_{n}$ is denoted by $\mathcal{R}_{1,n}$. In the routing problem, each vehicle $i$ is associated with a trip defined by its origin-destination (OD) pair, denoted as $(o_{i},d_{i})$. Namely, $v_{1}$ in (1) is initialized as $o_{i}$, and $v_{n}$ is fixed as $d_{i}$. The nodes $\{v_{k}\}$, $k\!=\!1,\dots,n\!-\!1$, in the feasible route are defined as a sequence of decision-making points (DMPs). At each DMP $v_{k}$, the vehicle’s route from $v_{k}$ to $d_{i}$ is optimized based on its location and real-time traffic conditions. This iterative process enables routing solutions to dynamically adapt to vehicle movements and evolving network dynamics.
II-B Vehicle Dynamics
To describe vehicle movements between DMPs, we introduce the following dynamic model. For each vehicle $i\!\in\!\mathcal{N}(t)$, let $a_{k}^{i}$ and $a_{k+1}^{i}$ be its arrival times at two consecutive DMPs $v_{k}$ and $v_{k+1}$, respectively. The dynamics of vehicle $i$ is then depicted as
$a^{i}_{k+1}=a^{i}_{k}+\tau_{k,k+1}^{i},$ (2)
where $\tau_{k,k+1}^{i}$ denotes the travel time of vehicle $i$ to traverse the edge $(v_{k},v_{k+1})$, which can be modeled using the widely adopted Bureau of Public Road (BPR) function [13]. The BPR model captures the impact of traffic congestion by describing how travel time increases as traffic demand approaches or exceeds road capacity. Specifically, $\tau_{k,k+1}^{i}$ can be denoted as
$\tau_{k,k+1}^{i}=\tau_{k,k+1}^{0}\Bigg{(}\!1\!+\!\alpha\!\left(\frac{f_{k,k+1}^{i}}{c_{k,k+1}}\right)^{\beta}\!\Bigg{)},$ (3)
where $\tau_{k,k+1}^{0}$ denotes the free-flow travel time on $(v_{k},v_{k+1})$, $f_{k,k+1}^{i}$ denotes the traffic flow on the edge when vehicle $i$ passes through, and $c_{k,k+1}$ is the road capacity. The parameters $\alpha$ and $\beta$ characterize the sensitivity of travel time to traffic flow, with typical values of $\alpha\!=\!0.15$ and $\beta\!=\!4$.
In a dynamic system, $f_{k,k+1}^{i}$ varies over time, depending on the number of vehicles traveling on $(v_{k},v_{k+1})$ within the same time interval. Next, we present appropriate monitoring and estimation approaches to characterize the traffic flow for each vehicle based on its arrival time at a DMP.
II-C Traffic Flow Monitoring &\&& Estimation
At a DMP $v_{k}$, the traffic flow on each edge forming the potential feasible routes from $v_{k}$ to $d_{i}$ needs to be monitored or estimated for optimal routing. To this end, we compute the traffic flow in two phases: (i) monitoring the traffic flow on adjacent edges, which are directly adjacent to $v_{k}$, and (ii) estimating the traffic flow on future edges, which are edges further along the feasible routes to the destination.
II-C1 Flow Monitoring for Adjacent Edges
Let $(v_{k},v_{k+1})$ be an adjacent edge along the feasible route $r_{k,d_{i}}$ from $v_{k}$ to $d_{i}$, where $v_{k+1}\!\in\!{\mathcal{V}_{k+1}}$ may not be unique with $\mathcal{V}_{k+1}$ denoting the set of alternative nodes for $v_{k+1}$. Given the arrival time $a_{k}^{i}$, a time window for monitoring the traffic flow on $(v_{k},v_{k+1})$ is defined as
$\mathcal{F}\big{(}a_{k}^{i},\Delta{t}\big{)}\!:=\!\big{[}a_{k}^{i}\!-\!\Delta{t},a_{k}^{i}\!+\!\Delta{t}\big{]},$ (4)
where $\Delta{t}\!>\!{0}$ defines the size of the monitoring window. Accounting for other vehicles $j\!\in\!\mathcal{N}(t)\!\setminus\!\{i\}$ entering the same edge within this window, we define the following indicator function
$\!\!\mathbf{1}_{a_{k,k+1}^{j}}\!:=\!\begin{cases}1,&\text{if~{} $a_{k,k+1}^{j}\!\in\!{\mathcal{F}\big{(}a_{k}^{i},\Delta{t}\big{)}}$,}\\ 0,&\text{otherwise,}\end{cases}$ (5)
where $a_{k,k+1}^{j}$ denotes the time when vehicle $j$ arrives at $v_{k}$ and enters $(v_{k},v_{k+1})$, satisfying $a_{k,k+1}^{j}\!=\!a_{k}^{j}$. As such, the traffic flow on each adjacent edge upon vehicle $i$’s arrival can be monitored and computed by
$\tilde{f}_{k,k+1}^{i}=\frac{\sum_{j\in\mathcal{N}(t)\setminus\{i\}}{\!\mathbf{1}_{a_{k,k+1}^{j}}\!+\!1}}{2\Delta{t}},$ (6)
including vehicle $i$ itself. The total length of the monitoring time window is $2\Delta{t}$, as defined in (4).
II-C2 Flow Estimation for Future Edges
For future edges $(v_{s},v_{s+1})\!\in\!{r_{k,d_{i}}}$, $s\!\neq\!{k}$, the traffic flow at a future arrival time $a_{s}^{i}$ cannot be directly observed at time $a_{k}^{i}$. However, by leveraging the route plans of other vehicles $j\!\in\!\mathcal{N}(t)\!\setminus\!\{i\}$, the traffic flow can be anticipated, enabling congestion estimation and route optimization for vehicle $i$ at each of its DMP $v_{k}$.
To proceed, let $r_{k^{\prime},d_{j}}^{*}(a_{k}^{i})$ be the optimal route plan of vehicle $j\!\in\!\mathcal{N}(t)\!\setminus\!\{i\}$ at time $a_{k}^{i}$ for completing its remaining trip, where $k^{\prime}$ indicates that the route plan was made at DMP $v_{k^{\prime}}$. For each future edge $(v_{s},v_{s+1})\!\in\!{r_{k,d_{i}}}$, $s\!\neq\!{k}$, the arrival time of vehicle $i$ at $v_{s}$ can be estimated by (2), i.e.,
$\hat{a}_{s}^{i}=\begin{cases}a_{k}^{i}+\tilde{\tau}_{k,k+1}^{i},&\text{if $s\!=\!k\!+\!1$,}\\ \hat{a}_{s-1}^{i}+\hat{\tau}_{s-1,s}^{i},&\text{otherwise,}\end{cases}$ (7)
where $\tilde{\tau}_{k,k+1}^{i}$ representing the time for vehicle $i$ to traverse edge $(v_{k},v_{k+1})$ is computed by integrating the monitored traffic flow $\tilde{f}_{k,k+1}^{i}$ in (3). Additionally, $\hat{\tau}_{s-1,s}^{i}$ in (7) denotes the estimated travel time for vehicle $i$ to traverse the future edge $(v_{s-1},v_{s})$. Using the estimated arrival time $\hat{a}_{s}^{i}$, the traffic flow on the future edge $(v_{s},v_{s+1})$ can be estimated by
$\hat{f}_{s,s+1}^{i}\!=\!\frac{\sum_{(v_{s},v_{s+1})\in{r}_{k^{\prime},d_{j}}^{*}\!(a_{k}^{i}),j\in\mathcal{N}(t)\setminus\{i\}}\!{\mathbf{1}_{\hat{a}_{s,s+1}^{j}}}\!\!+\!1}{2\Delta{t}},$ (8)
where $\hat{a}_{s,s+1}^{j}$ denotes the estimated time when vehicle $j$ arrives at $v_{s}$ and enters $(v_{s},v_{s+1})$, which is determined by the dynamic model of vehicle $j$ considering nominal travel times along $r_{k^{\prime},d_{j}}^{*}(a_{k}^{i})$. Similar to (5), the indicator function $\mathbf{1}_{\hat{a}_{s,s+1}^{j}}\!$ in (8) is defined as
$\mathbf{1}_{\hat{a}_{s,s+1}^{j}}\!:=\!\begin{cases}1,&\text{if~{}$\hat{a}_{s,s+1}^{j}\!\in\!{\mathcal{F}\big{(}\hat{a}_{s}^{i},\Delta{t}\big{)}}$,}\\ 0,&\text{otherwise.}\end{cases}$ (9)
By integrating $\hat{f}_{s,s+1}^{i}$ into (3), the estimated travel time $\hat{\tau}_{s,s+1}^{i}$ and the estimated arrival time $\hat{a}_{s+1}^{i}$ can be obtained accordingly. Note that for each future edge $(v_{s},v_{s+1})\!\in\!{r_{k,d_{i}}}$, $s\!\neq\!{k}$, the estimated traffic flow $\hat{f}_{s,s+1}^{i}$, travel time $\hat{\tau}_{s,s+1}^{i}$, and the resulting estimated arrival time $\hat{a}_{s+1}^{i}$ are computed iteratively.
II-D Routing Problem
In our exposition, we consider an emerging transportation system that integrates diverse transportation modes, including (1) private vehicles, (2) autonomous vehicles, and (3) ride-hailing vehicles. The sets of vehicles in each category are denoted by $\mathcal{N}_{p}(t)$, $\mathcal{N}_{a}(t)$, and $\mathcal{N}_{h}(t)$, respectively, with $\mathcal{N}(t)\!=\!\mathcal{N}_{p}(t)\cup{\mathcal{N}_{a}(t)}\cup{\mathcal{N}_{h}(t)}$. For simplicity, we assume travelers taking ride-hailing vehicles share the same OD pairs and evenly split the travel expenses. Private and autonomous vehicles, however, serve only one traveler at a time. Owing to the automation technology, the per-mile travel cost of autonomous vehicles is typically lower than that of private vehicles.
While individuals select their travel modes based on factors such as income, accessibility, and reliability, the quality of their trips should remain comparable to ensure equitable access to transportation resources and services. To this end, this paper aims to develop a D-RGS that assists each vehicle $i\!\in\!\mathcal{N}(t)$ in route planning by optimally selecting a route $r_{k,d_{i}}^{*}$ from $\mathcal{R}_{k,d_{i}}$ at each DMP $v_{k}$ while enhancing trip equity across all travelers.
III Dynamic Trip Index and Trip Equity
This section presents a framework to evaluate the quality of each trip and the equity of all trips in a dynamic transportation system. The framework is built upon two concepts: the dynamic trip index and dynamic trip equity that we introduce next.
III-A Dynamic Trip Index
The trip index quantifies the quality of a trip in terms of its efficiency (i.e., travel time), cost, and convenience, which are key factors in the evaluation of transport quality. In a dynamic transportation system, the trip index is influenced by factors such as the travel mode selected, the vehicle’s routing, and the schedules and routes of other vehicles. For any traveler taking a vehicle $i\!\in\!\mathcal{N}(t)$ for the trip $(o_{i},d_{i})$, the Dynamic Trip indeX (DTX) evaluated at time $t$ is defined as
$DTX_{i}(t):=\xi_{1}\frac{\tau_{\min}^{i}}{\tau_{i}(t)}+\xi_{2}\frac{\phi_{\min}^{i}}{\phi_{i}(t)}+\xi_{3}\frac{q_{\min}^{i}}{q_{i}},$ (10)
where the three terms represent the evaluations of efficiency, cost, and convenience of the trip, respectively. Specifically, $\xi_{1},\xi_{2},\xi_{3}\!\in\!\mathbb{R}_{+}$ are positive constants serving as weighting coefficients to balance the contribution of each factor in $DTX_{i}(t)$. These parameters satisfy the convex combination condition $\xi_{1}\!+\!\xi_{2}\!+\!\xi_{3}\!=\!1$, ensuring that the weights are normalized and appropriately distributed across the evaluated factors. In (10), $\tau_{\min}^{i}$, $\phi_{\min}^{i}$, and $q_{\min}^{i}$ denote the minimum values of travel time, cost, and inconvenience achievable for completing the trip $(o_{i},d_{i})$ among all alternative transportation modes under ideal traffic conditions (i.e., without congestion), which are computed by
$\tau_{\min}^{i}$ $\in\arg\min_{r\in\mathcal{R}_{o_{i},d_{i}}}\!\sum_{(v_{\ell},v_{\ell+1})\in{r}}\!\!\!\!\!\!\tau^{0}_{\ell,\ell+1},$ (11a)
$\phi_{\min}^{i}$ $=\epsilon_{\min}^{i}\tau_{\min}^{i},$ (11b)
$q_{\min}^{i}$ $=\frac{T_{w,\min}^{i}}{T_{d,\max}^{i}},$ (11c)
where, as defined earlier, $\mathcal{R}_{o_{i},d_{i}}$ is the set of all feasible routes from $o_{i}$ to $d_{i}$ and $\epsilon_{\min}^{i}\!:=\!\min\big{\{}\epsilon_{p},\epsilon_{a},\epsilon_{h}\big{\}}$ represents the minimum transportation cost per traveler per unit of travel time, with $\epsilon_{p}$, $\epsilon_{a}$, $\epsilon_{h}$ denoting the cost for private, autonomous, and ride-hailing vehicles, respectively. In (11c), the best convenience of the trip (i.e., minimum inconvenience) is denoted by $q_{\min}^{i}$, where
$T_{w,\min}^{i}:\!$ $=\min\big{\{}T_{w,p},T_{w,a},T_{w,h}\big{\}},$ (12a)
$T_{d,\max}^{i}:\!$ $=\max\big{\{}T_{d,p},T_{d,a},T_{d,h}\big{\}}.$ (12b)
We make use of $T_{w,p}$, $T_{w,a}$, $T_{w,h}$ to denote the average waiting times to access a private, autonomous, and ride-hailing vehicle, respectively. Here, $T_{w,p}$ accounts for the additional time required to park and retrieve a private vehicle, $T_{w,a}$ captures the potential delays due to scheduling and operational constraints of an autonomous vehicle, and $T_{w,h}$ reflects the average waiting time to dispatch a ride-hailing vehicle. In (12b), $T_{d,p}$, $T_{d,a}$, $T_{d,h}$ denote the feasible departure time windows (in hours per day) for each type of vehicle. It is worth noting that $T_{w,i}$ quantifies the accessibility of vehicle $i$, while $T_{d,i}$ characterizes its availability.
The dynamic values of trip time, cost, and convenience are denoted by $\tau_{i}(t)$, $\phi_{i}(t)$, and $q_{i}$ in (10), respectively. Let $r_{i}^{*}(t)$ be the optimal route plan for vehicle $i$ at time $t$ from its origin to the destination, which can be represented as
$r_{i}^{*}(t)=r_{o_{i},k}(t)\cup{r_{k,d_{i}}^{*}\!(t)},$ (13)
where $t\!\in\![a_{k}^{i},a_{k+1}^{i})$ and $r_{o_{i},k}(t)$ denotes the experienced route of vehicle $i$ at time $t$. The optimal route planning for completing the rest of the trip is denoted as $r_{k,d_{i}}^{*}(t)$, which remains unchanged until the vehicle reaches its next DMP $v_{k+1}$. In light of (13), we have
$\!\!\!\!\tau_{i}(t)$ $=\!\!\!\sum_{(v_{\ell},v_{\ell+1})\in{r_{o_{i},k}(t)}}\!\!\!\!\!\!\!\!\!\!\!\!\!\tau_{\ell,\ell+1}^{i}\!+\tilde{\tau}_{k,k+1}^{i}+\!\!\!\!\!\!\!\!\!\sum_{(v_{s},v_{s+1})\in{r_{k,d_{i}}^{*}\!(t)}}\!\!\!\!\!\!\!\!\!\!\!\!\!\hat{\tau}_{s,s+1}^{i},$ (14a)
$\phi_{i}(t)$ $=\epsilon_{i}\tau_{i}(t),$ (14b)
$q_{i}$ $=\frac{T_{w,i}}{T_{d,i}},$ (14c)
where $(v_{k},v_{k+1})\!\in\!{r_{k,d_{i}}^{*}\!(t)}$ and $s\neq{k}$ in (14a). The estimated travel times $\tilde{\tau}_{k,k+1}^{i}$ and $\hat{\tau}_{s,s+1}^{i}$ are computed using the monitored and estimated traffic flow given in (6) and (8), respectively. In addition, $\epsilon_{i}\!\in\!\{\epsilon_{p},\epsilon_{a},\epsilon_{h}\}$, $T_{w,i}\!\in\!\{T_{w,p},T_{w,a},T_{w,h}\}$ and $T_{d,i}\!\in\!\{T_{d,p},T_{d,a},T_{d,h}\}$, relying on the type of vehicle $i$.
Remark 1
The values of $\tau_{\min}^{i}$, $\phi_{\min}^{i}$, and $q_{\min}^{i}$ are achieved by selecting the most favorable vehicle type, which offers the maximum convenience, the shortest travel time, and operates under ideal traffic conditions without congestion. By comparing the actual travel conditions to this optimal scenario, the dynamic metric $DTX_{i}(t)$ quantifies the trip quality of the traveler in vehicle $i$ relative to the best possible travel experience.
III-B Dynamic Trip Equity
Dynamic Trip Equity (DTE) assesses how equitably the transportation resources and services are distributed among all travelers. In a dynamic transportation system, vehicles with overlapping edges in routes compete for road resources, leading to potential congestion. To capture the impact of competing vehicles on traffic conditions and routing decisions, we define the road resource competitors (RRCs) for vehicle $i\!\in\!\mathcal{N}(t)$ at time $t\!\in\![a_{k}^{i},a_{k+1}^{i})$ as
$\mathcal{C}_{i}(t)\!:=\!\big{\{}j\!\in\!\mathcal{N}(t)\big{|}r_{k^{\prime},d_{j}}^{*}\!(t)\cap{r_{k,d_{i}}}\!\neq\!{\emptyset},r_{k,d_{i}}\!\!\in\!{\mathcal{R}_{k,d_{i}}}\big{\}},\!$ (15)
where $r_{k^{\prime},d_{j}}^{*}\!(t)$ denotes the optimal route plan of vehicle $j$ at time $t$ to complete the rest of its trip. Vehicles whose routes do not share any overlapping edges with the feasible route of vehicle $i$ are not directly competing for the same road resources and, thus, are not included in $\mathcal{C}_{i}(t)$.
Instead of evaluating the DTE among all travelers in $\mathcal{N}(t)$, which can be computationally inefficient, we propose the DTE defined within the RRCs for each vehicle based on their route plans, ignoring indirect couplings between vehicles. Let $\mathcal{C}_{i,p}(t)$, $\mathcal{C}_{i,a}(t)$, $\mathcal{C}_{i,h}(t)$ be the sets of private, autonomous, and ride-hailing vehicles in $\mathcal{C}_{i}(t)$, i.e., $\mathcal{C}_{i}(t)\!=\!\mathcal{C}_{i,p}(t)\cup{\mathcal{C}_{i,a}(t)}\cup{\mathcal{C}_{i,h}(t)}$. We employ the Gini Coefficient [14], an effective measure of inequity in economics and social sciences, to quantify the DTX distribution among travelers. The Gini Coefficient is built upon the Lorenz curve [15], and by its definition, the DTE for the traveler in vehicle $i$ and those in the RRCs of vehicle $i$ at $t$ is defined as
$\!DTE_{i}(t)\!:=\!1\!-\!\frac{\sum_{j\in{\mathcal{S}_{i}(t)}}\!\sum_{j^{\prime}\in{\mathcal{S}_{i}(t)}}\!\big{|}DTX_{j}(t)\!-\!DTX_{j^{\prime}}(t)\big{|}}{2|\mathcal{S}_{i}(t)|^{2}{DTX_{i,\text{mean}}(t)}},$ (16)
where
$\mathcal{S}_{i}(t)=\big{(}\mathcal{C}_{i,p}(t)\cup\mathcal{C}_{i,a}(t)\big{)}\uplus\underbrace{\mathcal{C}_{i,h}(t)\dots\uplus\mathcal{C}_{i,h}(t)}_{\text{$m$}},$ (17a)
${DTX}_{i,\text{mean}}(t)=\frac{1}{|\mathcal{S}_{i}(t)|}\!\sum_{j\in{\mathcal{S}_{i}(t)}}\!\!\!\!DTX_{j}(t).$ (17b)
[…]
III-C Dynamic Trip Equity in Free-Flow Networks
This subsection discusses the DTE in a free-flow road network. In scenarios where the infrastructure is sufficient or the number of vehicles in the system is well below the road capacity, congestion is not a concern in route planning. In these cases, vehicles can freely choose the shortest routes for their trips. To describe the trip quality in this case, let $DTX_{i,0}$ be the DTX for vehicle $i\!\in\!\mathcal{N}(t)$ traversing a free-flow network. Since $\tau_{i}(t)\!=\!\tau_{\min}^{i}$, by (10), (11b), and (14b), we have
$DTX_{i,0}=\xi_{1}\!+\xi_{2}\frac{\epsilon_{\min}^{i}}{\epsilon_{i}}\!+\xi_{3}\frac{q_{\min}^{i}}{q_{i}},$ (18)
which is independent of time $t$. In line with the DTE presented above, the following result holds.
Proposition 1
For any vehicle $i\!\in\!\mathcal{N}(t)$ traveling in a free-flow road network on the trip $(o_{i},d_{i})$, the best DTE can be achieved between the traveler in vehicle $i$ and those in its RRCs, denoted as $\mathcal{C}_{i,0}(t)$, if the following condition is satisfied:
$DTX_{j,0}^{p}\!=\!DTX_{j,0}^{a}\!=\!DTX_{j,0}^{h},\quad\ \forall{j}\!\in\!\mathcal{C}_{i,0}(t),$ (19)
where
$DTX_{j,0}^{p}$ $=\xi_{1}\!+\xi_{2}\frac{\epsilon_{\min}^{j}}{\epsilon_{j}}\!+\xi_{3},\quad j\!\in\!\mathcal{C}_{i,0}^{p}(t),$ (20a)
$DTX_{j,0}^{a}$ $=\xi_{1}\!+\xi_{2}\!+\xi_{3}\frac{q_{\min}^{j}}{q_{j}},\quad j\!\in\!\mathcal{C}_{i,0}^{a}(t),$ (20b)
$DTX_{j,0}^{h}$ $=\xi_{1}\!+\xi_{2}\frac{\epsilon_{\min}^{j}}{\epsilon_{j}}\!+\xi_{3}\frac{q_{\min}^{j}}{q_{j}},\quad j\!\in\!\mathcal{C}_{i,0}^{h}(t).$ (20c)
Here, $\mathcal{C}_{i,0}^{p}(t)$, $\mathcal{C}_{i,0}^{a}(t)$, $\mathcal{C}_{i,0}^{h}(t)$ represent the sets of private, autonomous, and ride-hailing vehicles in $\mathcal{C}_{i,0}(t)$, respectively, with $\mathcal{C}_{i,0}(t)\!=\!\mathcal{C}_{i,0}^{p}\!\cup\!{\mathcal{C}_{i,0}^{a}(t)}\!\cup\!{\mathcal{C}_{i,0}^{h}(t)}$. Similar to (15), for $t\!\in\![a_{k}^{i},a_{k+1}^{i})$, $\mathcal{C}_{i,0}(t)$ is of the form
$\mathcal{C}_{i,0}(t)=\big{\{}j\!\in\!\mathcal{N}(t)\big{|}\ \underline{r}_{k^{\prime},d_{j}}\!(t)\cap{\underline{r}_{k,d_{i}}}\!\neq\!{\emptyset}\big{\}},$
where $\underline{r}_{k^{\prime},d_{j}}\!(t)$ denotes the shortest route for vehicle $j$ to complete its remaining trip and $\underline{r}_{k,d_{i}}$ is the shortest route from $v_{k}$ to $d_{i}$, considering free-flow travel times on edges.
[…]
Proposition 1 reveals the conditions required for maximizing trip equity across travelers in a free-flow road network. These conditions impose constraints on road infrastructure, transportation costs, and vehicle scheduling, which collectively determine the achievable DTX for each vehicle. Note that meeting these conditions ensures that the upper bound of the DTE attainable through route optimization is $1$. In other words, the D-RGS improves trip equity only through balancing the DTX via reducing traffic congestion. However, it cannot mitigate inequities caused by other factors, such as travel costs or limitations inherent to certain vehicle types.
IV Routing Guidance System Design with Improved DTE
This section proposes a D-RGS that optimally recommends vehicle routes to enhance $DTE_{i}(t)$ while considering potential traffic congestion. Let $d^{i}$ denote the departure time of vehicle $i\!\in\!\mathcal{N}(t)$ from its origin, i.e., $a_{o_{i}}^{i}\!\!=\!d^{i}$. For any vehicle $i$ arriving at a DMP $v_{k}$ at time $t\!=\!a_{k}^{i}$, the D-RGS provides an optimal route recommendation to the vehicle by addressing
$\max_{r_{k,d_{i}}\in{\mathcal{R}_{k,d_{i}}}}DTE_{i}(a_{k}^{i})$ (21)
subject to the constraints
$\hat{a}_{k+1}^{i}\!=\!a_{k}^{i}\!+\tilde{\tau}_{k,k+1}^{i},\ \ (v_{k},\!v_{k+1})\!\in\!{r_{k,d_{i}}},\!\!$ (22a)
$\hat{a}_{s+1}^{i}\!=\!\hat{a}_{s}^{i}\!+\hat{\tau}_{s,s+1}^{i},\ \ (v_{s},\!v_{s+1})\!\in\!{r_{k,d_{i}}},\ s\!\neq\!{k},$ (22b)
$a_{k,k+1}^{j}\!=\!a_{k}^{j},\ \ j\!\in\!\mathcal{C}_{i}(a_{k}^{i}),\ j\!\neq\!{i},$ (22c)
$\hat{a}_{h,h+1}^{j}\!=\!\hat{a}_{h}^{j},\ j\!\in\!\mathcal{C}_{i}(a_{k}^{i}),\ j\!\neq\!{i},\ h\!\neq\!{k},$ (22d)
$\hat{a}_{h+1}^{j}\!=\!\hat{a}_{h}^{j}\!+\!\tau^{0}_{h,h+1},\ \ (v_{h},\!v_{h+1})\!\in\!{r_{k^{\prime},d_{j}}^{*}\!(a_{k}^{i})},$ (22e)
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ j\!\in\!\mathcal{C}_{i}(a_{k}^{i}),\ j\!\neq\!{i},$
[…]
In the optimization problem formulated above, the target is to maximize $DTE_{i}(t)$, as defined in (15)-(17), by optimally selecting a route $r_{k,d_{i}}$ from the set of all feasible routes $\mathcal{R}_{k,d_{i}}$. Constraints (22a) and (22b) provide estimated dynamics for vehicle $i$, while (22c)-(22e) describe the estimated dynamics for vehicles $j\!\in\!\mathcal{C}_{i}(a_{k}^{i})$, $j\!\neq\!{i}$. Travel times $\tilde{\tau}_{k,k+1}^{i}$ and $\hat{\tau}_{s,s+1}^{i}$, derived from the monitored and estimated traffic flow, are computed by (3)-(6), (8) and (9). The computation of the DTX is based on the constraints (10)-(14).
The problem (21) subject to (22) can be addressed by dynamic programming [16, 17] each time a vehicle reaches a DMP, and the optimal solution is denoted as $r_{k,d_{i}}^{*}$. It is worth noting that, if vehicle $i$ has no RRCs at $v_{k}$, i.e., $|\mathcal{C}_{i}(a_{k}^{i})|\!=\!1$, the D-RGS then provides the shortest route for vehicle $i$, as $DTE_{i}(a_{k}^{i})\!=\!1$ naturally holds in this case. Unlike one-time optimization, the D-RGS continuously re-optimizes and updates the route recommendation at each MDP by maximizing the dynamic trip equity $DTE_{i}(a_{k}^{i})$. This facilicates incorporating real-time traffic data and coping with environmental uncertainties.
Remark 2
The computational complexity for solving (21) subject to (22) is determined by $|\mathcal{R}_{k,d_{i}}|$ and $|\mathcal{C}_{i}(a_{k}^{i})|$. To achieve a trade-off between the control performance and the computational efficiency, Yen’s algorithm [18] is employed to construct $\mathcal{R}_{k,d_{i}}$ at each DMP, which selects the $L$ shortest routes from $v_{k}$ to $d_{i}$ instead of using all feasible routes. Notice that both $|\mathcal{R}_{k,d_{i}}|$ and $|\mathcal{C}_{i}(a_{k}^{i})|$ decrease as the vehicle approaches its destination, which further enhances the computational efficiency of our method.
Remark 3
The D-RGS provides socially optimal route recommendations to individual vehicles that enhance trip equity across travelers. In practice, vehicles may deviate from the recommended routes based on their own route optimization metric or preference [19], such as minimizing travel time or selecting routes with better road conditions. Our follow-up work [20] focuses on effective incentive design to improve route compliance and social optimality.
V Simulation Studies
This section performs simulation studies to illustrate the effectiveness of the proposed framework and the D-RGS.
V-A Setup
Figure 1: An urban road network in Boston.
As shown in Fig. 1, we consider an urban road network in Boston that comprises $58$ nodes, including $45$ road intersections (in blue), $8$ origins (in green), and $5$ destinations (in red). The simulation includes $1,000$ vehicles with random departure times between $08\!\!:\!\!00$-$10\!\!:\!\!00$. The OD pair of each vehicle is randomly selected from the origin and destination sets. Based on indices, vehicles are categorized as: private vehicles ($1\!-\!500$), autonomous vehicles ($501\!-\!800$), and ride-hailing vehicles ($801\!-\!1000$), where each ride-hailing vehicle is shared by two travelers, i.e., $m\!=\!2$ in (17a). All vehicles travel at a constant free-flow speed of $27$ km/h, and the nominal travel time on each edge is obtained from OpenStreetMap [21]. The maximum traffic flow on each edge is considered as $5$ vehicles per minute per lane. The parameter settings for each vehicle type satisfying the conditions in Proposition 1 are provided in Table I. To enhance the computational efficiency, Yen’s algorithm is employed to select $7$ shortest routes from a DMP $v_{k}$ to the destination $d_{i}$ when forming the feasible route set $\mathcal{R}_{k,d_{i}}$.
TABLE I: Parameter Settings of Each Vehicle Type
Vehicle type $\xi_{1}$ $\xi_{2}$ $\xi_{3}$ $\epsilon_{i}$[\$/min] $T_{w,i}$[min] $T_{d,i}$[h]
$i\!\in\!\mathcal{N}_{p}(t)$ $0.4$ $0.4$ $0.2$ $0.27$ $2$ $24$
$i\!\in\!\mathcal{N}_{a}(t)$ $0.4$ $0.4$ $0.2$ $0.1485$ $15$ $18$
$i\!\in\!\mathcal{N}_{h}(t)$ $0.4$ $0.4$ $0.2$ $0.1536$ $6$ $12$
Figure 2: Comparison of the DTX. The results show that DTXs in our method cluster more closely around the average value.
V-B Results and Analysis
We compare the proposed method with the pre-planned and dynamic shortest-route strategies, where the pre-planned shortest-route strategy minimizes the trip time based on the nominal travel time on each edge, while the dynamic strategy minimizes the travel time dynamically using the monitored and estimated traffic flow information. Traffic congestion is monitored in real-time at intervals of $\Delta{t}\!=\!60$ seconds. The DTX of each vehicle integrating the actual travel time in (10) in the three strategies is shown in Fig. 2. As shown in Fig. 2(a), the pre-planned strategy results in a wide distribution of trips with very high and low dynamic trip indices, while trips in the mid-range are relatively sparse. The dynamic shortest-route strategy (see Fig. 2(b)), which incorporates real-time traffic information and improves performance for some vehicles with higher DTX, yet a significant fraction of vehicles still experiences a small DTX value. In contrast, the proposed method leads to a more balanced distribution, as shown in Fig. 2(c), where trip indices converge closer to the average level. This improvement enhances the DTE for all trips from $0.734$ to $0.818$, demonstrating the effectiveness of our approach. Notably, the DTE here is computed by
$DTE=1-\frac{\sum_{i=1}^{S}\sum_{i^{\prime}=1}^{S}\!|DTX_{i}\!-\!DTX_{i^{\prime}}|}{2S^{2}DTX_{\text{mean}}}.$
Similar to (16), $S\!=\!1200$ denotes the total number of traveler trips. Note that each ride-hailing vehicle serves two travelers. Thus, $DTX_{i+200}\!=\!DTX_{i}$ for $i\!=\!801,\dots,1000$.
Figure 3: Comparison of the average trip time.
Figure 4: Comparison of the average trip cost.
The comparison of the average trip time and cost in the three methods are shown in Fig. 3 and Fig. 4, respectively. We observe that, in comparison with the pre-planned shortest-route strategy, our method significantly reduces the average trip time for private, ride-hailing, and autonomous vehicles by approximately $30\%$, $33\%$, and $42\%$, respectively, leading to savings of about \$$\text{\$}5.7$, \$$\text{\$}2.7$, and \$$\text{\$}3.6$ per trip for each type of vehicle. It is also interesting to see that, compared to the dynamic shortest-route strategy, our method slightly increases the average trip time for private vehicles by $3.1$ minutes, leading to \$$\text{\$}0.84$ increase in the average trip cost. However, it contributes to reducing the trip time and cost for both ride-hailing and autonomous vehicles by approximately $8.8\%$ and $14.7\%$, respectively. This aligns with real-world expectations, as private vehicle owners typically belong to higher-income groups, whereas individuals with lower incomes are more likely to take cost-effective travel options such as ride-hailing and autonomous vehicles. The simulation results above show that our method redistributes travel costs across different vehicle types through route optimization, contributing to a more equitable and efficient transportation system.
VI Conclusion
In this paper, we presented a framework to quantify trip quality and equity in a dynamic emerging transportation system, integrating trip time, cost, and convenience. We developed a routing guidance system to optimize vehicle route recommendations, incorporating real-time and anticipated traffic congestion while improving trip equity. In addition, we established conditions that ensured perfect trip equity in free-flow networks. Our simulation studies on a Boston road network demonstrated that the proposed approach increased trip equity from $0.734$ to $0.818$ compared to the pre-planned shortest-route strategy, leading to $11.4\%$ improvement and a more equitable transportation system. Future work will focus on learning compliance probability and designing effective incentive mechanisms to enhance drivers’ alignment with routing recommendations, thereby improving social equity and system-wide optimality.